Present: R. Ostojic, M. Fraser, J. Jentzsch, W. Bartmann
Elena TL status:
- Optics layout being finalized, document being prepared
- LNS commissioning: diff pumping stage available at end of the year, SEM not known when, would need some money for dummy chambers, detectors, supports, cables
- Need clear statement from Christian and Tommy if LNS commissioning prior to machine commissioning is of interest
- Quadrupole NEG coating not performing as expected; first bakeout below NEG activation temp OK, but NEG activation requires higher temperature than expected and does not pump as expected; chemical cleaning seems OK; all done within VSC;
- One quad assembly will be dissassembled to check pumping of only the chamber (Sept-15?, VSC)
- Discussions about TL vacuum requirements ongoing
- Add to discussion with experiments (1-October)
- Prepare TL vacuum specification document (1-October)
- Prepare baseline of experiment handover positions, optical and physical (1-Oct)
- Status of deflectors:
- First unit 33 deg (2 subfamilies) finished (little modifs this pm)
- Second unit will be a lot of copy of first unit; relatively straightforward
- Third unit (vertical bend of 80 deg) to be started
- Status of quads:
- Batch of 12 units ready for NEG coating, but wait for results of opened assembly
- Rest will be delivered at end of month
- Magnetic shielding
- Incorporate stray field measurements in beam optics; simulate solenoid field acc to measurements and translate into multipole kicks on beam; time varying; take into account all possible errors (field, misalignment)
- Prepare correction procedure as operational tool for commissioning (March-16)
- Phase 2/3 installation: timing not clear, manpower planning
There are minutes attached to this event.
Show them.